How can $58,000 be invested, part at 10% annual simple interest and the remainder at 8% annual simple interest, so that the interest earned by the two accounts will be equal.
Please help. I can set up the equation if the total annual simple interest earned for 1 year was given, but this one is a little over my head.

Answer by checkley77(12844)

Set up an equation with the unknown amount (x) @ 10% = (58,000-x) amount @ 8%.
.10x=.08(58,000-x)
.10x=4,640-.08x
.10x+.08x=4,640
.18x=4,640
x=4,640/.18
x=25,777.78 amount invested @ 10%.
58,000-25,777.77=32,222.23 amount invested @ 8%.
Proof:
.10*25,777.77=.08*32,222.23
2,577.77=2,577.77